Weight of oxygen in 1kg of water is:
A. 800gm
B. 889gm
C. 112gm
D. None

Grade:12th pass

2 Answers

Rajat
4 years ago
Water is H2O 
Molar mass of H2O is 18 gm and mass of oxygen in 1 mole H2O is 16 gm.
So in 1 kg H2O mass of Oxygen is : 1000×16/18 gm = 888.889= 889 gm
Arun
4 years ago
molar mass H20 =- 18g/mol  
mol H2O in 1000g H2O = 1000/18 = 55.56mol  
Water : 1mol H2 = 2.0g/mol plus 1 mol O = 16g/mol  
Mass of H2 in 55.56 mol H2O = 55.56*2.0 = 111.1g  
Mass of O in 55.566 mol O = 55.56*16 =888.9 g  
1kg H2O contains 111.1g hydrogen and 888.9g oxygen